# HydroGEM: A Self-Supervised Zero Shot Hybrid TCN-Transformer Foundation Model for Continental Scale Streamflow Quality Control

### v2

@@ 6.2 Limitations and Operational Considerations
Despite strong performance, several limitations require careful consideration for operational deployment. These constraints arise from both technical architecture decisions and fundamental challenges in defining “correct” hydrological data.
@@
Reconstruction divergence in ice-affected periods represents the most significant operational challenge. While HydroGEM successfully detects ice-related anomalies with high recall, reconstructions during these periods often diverge from operational corrections, particularly for Canadian stations. This divergence reflects fundamental differences in correction philosophy between agencies. USGS guidance emphasizes maintaining physical plausibility of both discharge and stage, typically applying moderate adjustments that preserve the coupled relationship. In contrast, ECCC practice often aggressively reduces discharge to near-zero while preserving elevated stage readings during ice backwater, acknowledging that stage may accurately represent local conditions even when unsuitable for discharge computation [18]. This is not a detection failure but rather highlights that “correct” values depend on operational context, downstream use cases, and agency-specific protocols. For deployment, we strongly recommend treating all reconstructions as suggestions requiring expert review, particularly during known ice periods.
@@
The model occasionally flags rapid but valid discharge fluctuations from gate operations, hydropeaking, or flash floods as anomalous. These events produce sudden excursions in discharge-stage relationships that resemble sensor artifacts in the learned representation. While maintaining high recall ensures genuine anomalies are rarely missed (a critical requirement for operational systems), this behavior reinforces the necessity of human oversight. The challenge of distinguishing anthropogenic flow modifications from sensor failures without explicit metadata represents a fundamental limitation of approaches based solely on time series patterns. Future work could incorporate auxiliary information about control structures, precipitation events, or operational schedules to improve discrimination.
@@
Training requires substantial computational resources, with backbone pretraining consuming approximately 48 GPU-hours on modern hardware. While this one-time cost amortizes across thousands of deployment sites, it may limit accessibility for smaller agencies or research groups. Inference remains efficient at less than 100ms per window on standard hardware, enabling real-time deployment. However, the current architecture processes 576-hour windows in batch mode, creating a tradeoff between temporal context and streaming responsiveness. Operational systems requiring immediate anomaly detection would benefit from architectural modifications enabling incremental updates while maintaining long-term memory.
@@
The model operates solely on observable quantities without meteorological forcing, maximizing applicability to gauge-only stations but preventing discrimination between sensor failures and physical extremes using precipitation context. A 500% discharge spike could represent sensor malfunction or an actual flash flood; without rainfall data, the model cannot definitively distinguish these cases. Integration with gridded meteorological products or real-time weather station networks could improve discrimination, though at the cost of increased system complexity and reduced deployability to remote sites lacking complementary observations.
6.3 Advancing the Field
HydroGEM advances hydrological quality control beyond previous approaches along multiple dimensions. Traditional rule-based systems require extensive manual calibration and capture only simple threshold violations. Even the strongest baseline overall, Isolation Forest, achieves only F1 = 0.392 on synthetic data despite being a general-purpose unsupervised anomaly detector. Moreover, baseline performance varies dramatically between evaluation settings: Isolation Forest drops to F1 = 0.263 on Canadian data, while Persistence rises from F1 = 0.001 (synthetic) to F1 = 0.418 (Canadian), highlighting how specialized methods fail when anomaly distributions shift. Rating curve residual methods, despite their theoretical grounding in hydraulic principles, achieve F1 = 0.056 on synthetic data because they examine only instantaneous relationships while missing temporal evolution patterns. HydroGEM maintains strong performance across both evaluation settings through learned representations that adapt to local context while capturing both instantaneous and temporal dependencies.
Prior machine learning approaches in hydrology have typically focused on individual sites or small regional clusters, requiring site-specific training that limits scalability. Studies reporting strong performance often evaluate on single watersheds or validate using random train-test splits that allow information leakage between partitions. By training on 3,724 sites and evaluating on 799 completely unseen locations with rigorous geographic separation, we establish that multi-site learning with appropriate normalization achieves better generalization than specialized models. The zero-shot transfer to Canadian stations provides additional evidence that continental-scale training produces robust representations applicable beyond the training distribution.
Within the broader foundation model framework, HydroGEM represents the first systematic application to in situ environmental sensor networks. Weather and climate foundation models have achieved remarkable success with gridded reanalysis data, but sensor networks present unique challenges including irregular spatial coverage, extreme magnitude heterogeneity, sparse labeling, and mandatory human oversight requirements. Our two-stage training framework and hierarchical normalization scheme provide a template for adapting foundation model principles to these operational constraints. The approach could generalize to other environmental monitoring applications including water quality sensors, air quality networks, and ecological observation systems.
6.4 Broader Implications for Operational Hydrology
The development of HydroGEM has implications extending beyond immediate quality control applications. As monitoring networks expand globally, driven by decreasing sensor costs and climate adaptation needs, the gap between data generation and quality assurance capacity will continue widening. Current estimates suggest the global water monitoring network generates over 100 million observations daily, yet most water agencies employ fewer than 50 hydrologists for quality control. HydroGEM offers a path to scale expert oversight by automating routine screening while preserving human judgment for complex decisions. The three-tier flagging system enables efficient triage, passing high-confidence clean data with minimal review, prioritizing uncertain observations for inspection, and providing suggested corrections for detected anomalies.
The human-AI collaboration approach embodied by HydroGEM may serve as a model for AI integration in safety-critical environmental monitoring. Rather than pursuing full automation, the system amplifies human expertise through scalable pattern recognition while preserving accountability, interpretability, and corrective capacity. All outputs include provenance information suitable for regulatory audit. This design philosophy acknowledges that data quality decisions often involve value judgments about intended use, acceptable uncertainty, and operational constraints that extend beyond pure statistical pattern recognition.
Climate change intensifies the importance of reliable hydrological observations for adaptation planning, infrastructure design, and ecosystem management. As extreme events become more frequent and hydrological regimes shift beyond historical baselines, distinguishing sensor anomalies from genuine extremes becomes increasingly critical. HydroGEM’s ability to detect subtle violations of learned patterns while adapting to local context could help identify both data quality issues and emerging hydrological signatures of environmental change. The pretrained backbone could be fine-tuned for specific climate impact detection tasks, potentially serving as an early warning system for regime shifts.
@@ 7 Conclusion
This work introduced HydroGEM, a foundation model for continental-scale streamflow quality control that addresses the growing mismatch between automated data generation and expert review capacity in hydrological monitoring networks. Through careful architectural design and training strategies, we showed that foundation model principles can be successfully adapted to operational sensor networks despite challenges including extreme heterogeneity, sparse labeling, and stringent deployment constraints.
The technical contributions of this work center on three innovations. First, the two-stage training approach combines self-supervised pretraining on massive unlabeled data with synthetic anomaly fine-tuning, eliminating dependence on scarce labeled datasets while learning robust representations of hydrological normality. Second, the hierarchical normalization scheme allows learning across six orders of magnitude within a single model, addressing a key challenge in continental-scale hydrology. Third, the rigorous evaluation framework with expert-validated synthetic anomalies and zero-shot cross-national transfer provides strong evidence for genuine generalization rather than pattern memorization.
Experimental validation confirms the effectiveness of these approaches. On synthetic test data with known ground truth, HydroGEM achieves F1 = 0.792 with 68.7% reconstruction error reduction, outperforming the strongest baseline by 36.3% while maintaining 97% preservation of clean data. Zero-shot transfer to 100 Canadian stations yields F1 = 0.586, with detection patterns aligning closely with operational correction rates across seasons. The model successfully identifies all 18 expert-validated anomaly types and maintains consistent performance across correction magnitudes from 1% to 100%, indicating pattern-based rather than threshold-based detection.
Several avenues for future development emerge from this work. Immediate priorities include incorporating meteorological forcing to provide physical context for extreme values, developing streaming architectures for real-time deployment with incremental updates, and creating specialized fine-tuning procedures for ice-dominated watersheds where correction practices vary significantly between agencies. The modular architecture enables these extensions without complete retraining. Longer-term opportunities include multi-modal fusion with satellite observations for comprehensive watershed monitoring, extension to water quality parameters where sensor drift and biofouling present similar challenges, and development of active learning frameworks that continuously improve from operational corrections.
As environmental monitoring transitions from sparse manual sampling to dense automated sensing, quality control emerges as the critical bottleneck preventing full utilization of these data streams for scientific understanding and operational decision-making. HydroGEM shows that foundation models can help bridge this gap through human-AI collaboration that preserves expert oversight while enabling continental-scale monitoring. This approach, where AI handles pattern recognition at scale while humans provide domain expertise and accountability, may prove essential as we build the observational infrastructure needed to understand and manage water resources under accelerating environmental change.
@@ Acknowledgments
This material is based upon work supported by the National Science Foundation under Grant No. EAR 2012123. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation. Any use of trade, firm, or product names is for descriptive purposes only and does not imply endorsement by the U.S. Government. The work was also supported by the University of Vermont College of Engineering and Mathematical Sciences through the REU program. The authors would like to thank the US Geological Survey (USGS) for offering the domain expertise that was crucial to identify the peak anomaly types that are of practical importance. *** Ijaz, I think we should acknowledge USGS and ECCC for the datasets made available for us. *** The authors acknowledge the National Artificial Intelligence Research Resource (NAIRR) Pilot and the Texas Advanced Computing Center for contributing to the research by providing GPU computing resources under Award No. NAIRR240491.
@@
USGS streamflow data are publicly available through the National Water Information System (NWIS) at https://waterdata.usgs.gov/nwis. Environment and Climate Change Canada hydrometric data are available at https://wateroffice.ec.gc.ca/. Processed datasets and code will be made available upon publication at [repository URL to be added].
@@
Ijaz Ul Haq: Conceptualization, Methodology, Software, Validation, Formal analysis, Investigation, Data curation, Writing - Original Draft, Visualization. Byung Suk Lee: Supervision, Resources, Writing - Review & Editing, Funding acquisition. Julia N. Perdrial: Supervision, Domain expertise, Writing - Review & Editing, Funding acquisition.
@@ A.0.2 Tier 2: Site-Specific Standardization
$\mu_{\text{global}}=\mathbb{E}_{s\in\text{train}}\mathbb{E}_{t\in\text{train}(s)}[\mathbf{x}_{t}^{(1)}],\quad\sigma_{\text{global}}=\text{Std}_{s\in\text{train}}\text{Std}_{t\in\text{train}(s)}[\mathbf{x}_{t}^{(1)}]$ (9)
@@ A.0.5 Scale Embeddings
Design choice validation: Ablation studies (Section 5.4) show that removing scale embeddings degrades performance by 12-15% F1 score, confirming their necessity for continental-scale learning.
@@ C.1 Pretraining Loss
C.2 Fine-Tuning Loss
@@
Combined fine-tuning objective:
$\mathcal{L}_{\text{finetune}}=1.5\mathcal{L}_{\text{focal}}+\lambda_{c}(\epsilon)\mathcal{L}_{\text{corrupt}}+\lambda_{p}(\epsilon)\mathcal{L}_{\text{preserve}}+0.1\mathcal{L}_{\text{physics}}$ (39)
@@ D.0.4 Gap Statistics
  • 0% contain >>6 hours (sequences intersecting gaps >>6h are dropped)
Leakage prevention: Any sequence window intersecting a flagged timestep (outlier, physical implausibility, or gap >>6h) is excluded entirely, even if only a single hour is affected. This conservative approach eliminates potential information leakage from imputation.
@@ E.0.2 Coverage Control Algorithm
Appendix F Test-Time Anomaly Injection
This appendix provides complete mathematical formulations for the 18 test anomaly types, each with 3-4 equation variants.
F.0.1 Example: Drift Anomaly (4 variants)
Variant 1 - Linear:
@@ F.0.1 Example: Drift Anomaly (4 variants)
Variant 4 - Polynomial:
$Q^{\prime}(t)=Q(t)+[a_{Q}(t-t_{\text{start}})^{2}+b_{Q}(t-t_{\text{start}})],\quad\text{similarly for }H$ (48)
where coefficients $a,b$ chosen to achieve target endpoint deviation $\sim U(0.1Q_{\text{mean}},0.3Q_{\text{mean}})$
F.0.2 Example: Ice Backwater (3 variants)
Variant 1 - Gradual onset:
@@ F.0.2 Example: Ice Backwater (3 variants)
Variant 2 - Abrupt onset, gradual recovery:
@@
Variant 3 - Periodic breakup events:
@@
where $\alpha_{\text{base}}\sim U(0.2,0.4)$, $N_{\text{events}}\sim U(2,5)$, event amplitudes $A_{i}\sim U(-0.3,-0.1)$ (negative for breakup), widths $\sigma_{i}\sim U(1,6)$ hours
Appendix G Expert Review Documentation
This appendix documents the structured expert review process for synthetic anomaly validation.
G.0.1 Review Protocol
Domain experts evaluated each anomaly type using standardized forms. Reviewers included:
  • A faculty member in geoscience with extensive experience in water resources and hydrological research
  • A research assistant from the University of Vermont Water Resources Institute with operational experience from a USGS internship, including exposure to real-world streamflow data quality challenges
Realism assessment (4-point scale):
  • 4 - Very realistic: Commonly observed in operational QA/QC workflows
  • 3 - Realistic: Plausible, occasionally observed
  • 2 - Questionable: Theoretically possible but rarely encountered
  • 1 - Unrealistic: Does not match field experience
Temporal scale validation: For each anomaly type, reviewers identified duration ranges appearing implausible relative to operational experience. For example, ice backwater durations <<24 hours were flagged as too brief for realistic ice formation/breakup dynamics.
Confound detection: Reviewers flagged injection regimes creating:
  • Artificially easy detection (e.g., spike magnitudes $10\sigma$ never observed in practice)
  • Ambiguous labeling (e.g., patterns indistinguishable from natural variability)
  • Physical impossibilities (e.g., rating curve violations)
G.0.2 Key Feedback Integration
Table 9 documents major revisions based on expert feedback.
Table 9: Expert feedback integration
Topic Initial Expert Feedback Change Made
Ice duration Macro: 72-144h Ice persists weeks-months in northern systems Extended to 72-520h
Backwater onset 50% instantaneous Backwater develops gradually as obstruction forms All use gradual onset (>>6h)
Spike magnitudes Some $8\sigma$ Never see such extremes in approved data Capped at $5\sigma$
Coverage tier 40% probability Reduce for challenging evaluation Reduced to 20%
Equation variants Some had 2 Need stronger separation for generalization claims Minimum 3 per type
Drift subtlety All >>0.1 ft3s-1day-1 Drift can be much more subtle Added subtle variant
G.0.3 Consensus Categorization
The final 18-type taxonomy reflects reviewer agreement on field-relevance:
High confidence (9 types): All reviewers rated 3-4 (realistic to very realistic), with specific operational examples provided:
  • Dropout: Battery failure, telemetry loss
  • Flatline: Frozen float, clogged intake
  • Ice backwater: Common Nov-Apr northern stations
  • Rating drift: Frequent after major floods, channel migration
Medium confidence (4 types): Mixed ratings (2-4), considered plausible but less frequently documented:
  • Time misalignment: Possible with manual data entry, rare in modern systems
  • Gate operation: Relevant for regulated sites, not universal
Low confidence (4 types): Majority rated 2 (questionable), retained for completeness:
  • Splice: Hard to imagine operational scenario
  • Warp: Never seen systematic time stretching
Debated (retained with caveat): Reviewers disagreed on anomaly vs. natural variability:
  • Noise burst: Could be natural turbulence or sensor issue, ambiguous

### v3

@@ 6.2 Limitations and Operational Considerations
Despite strong performance, several limitations require careful consideration for operational deployment. These constraints arise from both technical architecture decisions and fundamental challenges in defining ground truth for hydrological data quality.
@@
Reconstruction divergence in ice-affected periods represents the most significant operational challenge. While HydroGEM successfully detects ice-related anomalies with high recall, reconstructions during these periods often diverge from operational corrections, particularly for Canadian stations. This divergence reflects fundamental differences in correction philosophy between agencies. USGS guidance emphasizes maintaining physical plausibility of both discharge and stage, typically applying moderate adjustments that preserve the coupled relationship. In contrast, ECCC practice often reduces discharge substantially while preserving elevated stage readings during ice backwater, acknowledging that stage may accurately represent local conditions even when unsuitable for discharge computation [20]. This is not a detection failure but rather highlights that appropriate corrections depend on operational context, downstream use cases, and agency-specific protocols. For deployment, we recommend treating all reconstructions as suggestions requiring review, particularly during known ice periods.
@@
The model occasionally flags rapid but valid discharge fluctuations from gate operations, hydropeaking, or flash floods as anomalous. These events produce sudden excursions in discharge-stage relationships that resemble sensor artifacts in the learned representation. While maintaining high recall ensures genuine anomalies are rarely missed, this behavior reinforces the necessity of human oversight. The challenge of distinguishing anthropogenic flow modifications from sensor failures without explicit metadata represents a fundamental limitation of approaches based solely on time series patterns.
@@
Training requires substantial computational resources, with backbone pretraining consuming approximately 48 GPU-hours on modern hardware. While this one-time cost amortizes across thousands of deployment sites, it may limit accessibility for smaller agencies or research groups. Inference remains efficient at less than 100ms per window on standard hardware, enabling real-time deployment. However, the current architecture processes 576-hour windows in batch mode, creating a tradeoff between temporal context and streaming responsiveness.
@@
The model operates solely on observable quantities without meteorological forcing, maximizing applicability to gauge-only stations but preventing discrimination between sensor failures and physical extremes using precipitation context. A large discharge spike could represent sensor malfunction or an actual flash flood; without rainfall data, the model cannot definitively distinguish these cases.
6.3 Future Directions
Several extensions could address current limitations and expand applicability. Integration of multivariate causal information represents a natural next step. Incorporating precipitation time series, upstream gauge observations, and reservoir release schedules would enable the model to reason about physical causes of observed patterns. For example, a discharge spike coinciding with heavy rainfall across the watershed likely reflects genuine hydrological response rather than sensor malfunction, while an isolated spike without precipitation or upstream response suggests instrumentation issues. Such causal reasoning could substantially reduce false positives on legitimate extreme events.
Development of streaming architectures would enable real-time deployment with incremental updates rather than batch processing. Architectures supporting online inference while maintaining long-term memory could reduce detection latency from hours to minutes, better matching operational monitoring requirements.
Integration with large language models offers potential for automated report generation. Detected anomalies could be summarized in natural language explanations describing the likely failure mode, affected time periods, and recommended actions, reducing cognitive load on operators reviewing flagged observations.
Extension to water quality parameters presents a natural application domain where sensor drift, biofouling, and calibration issues produce similar quality control challenges. The two-stage training framework and hierarchical normalization could transfer to dissolved oxygen, turbidity, conductivity, and other continuously monitored water quality variables.
Finally, active learning frameworks that continuously improve from operator correction feedback could enable ongoing model refinement as deployment generates new labeled examples, progressively improving discrimination in challenging cases like ice effects and controlled releases.
@@ 7 Conclusion
This work introduced HydroGEM, a foundation model for continental-scale streamflow quality control that addresses the growing mismatch between automated data generation and review capacity in hydrological monitoring networks. Through careful architectural design and training strategies, we demonstrated that foundation model principles can be successfully adapted to operational sensor networks despite challenges including extreme heterogeneity, sparse labeling, and stringent deployment constraints.
The technical contributions center on three innovations. First, the two-stage training approach combines self-supervised pretraining on 6.03 million sequences with Stage 2 training using synthetic anomalies, eliminating dependence on scarce labeled datasets while learning robust representations of hydrological normality. Second, the hierarchical normalization scheme enables learning across six orders of magnitude within a single model, addressing a key challenge in continental-scale hydrology. Third, the rigorous evaluation framework with synthetic anomalies and zero-shot cross-national transfer provides evidence for genuine generalization rather than pattern memorization.
[…]
As environmental monitoring transitions from sparse manual sampling to dense automated sensing, quality control emerges as the critical bottleneck preventing full utilization of these data streams for scientific understanding and operational decision-making. HydroGEM demonstrates that foundation models can help bridge this gap through human-AI collaboration that preserves oversight while enabling continental-scale monitoring. This approach, where AI handles pattern recognition at scale while humans provide domain expertise and accountability, may prove essential as we build the observational infrastructure needed to understand and manage water resources under accelerating environmental change.
Declaration of Generative AI and AI-Assisted Technologies in the Writing Process
During the preparation of this manuscript, the authors used ChatGPT (OpenAI) to assist with grammar and language refinement in the Introduction and Discussion sections. After using this tool, the authors reviewed and edited all content as needed and take full responsibility for the content of the publication.
@@ Acknowledgments
This material is based upon work supported by the National Science Foundation under Grant No. EAR 2012123. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation. Any use of trade, firm, or product names is for descriptive purposes only and does not imply endorsement by the U.S. Government.
The work was also supported by the University of Vermont College of Engineering and Mathematical Sciences through the REU program.
The authors acknowledge the U.S. Geological Survey (USGS) for providing access to the National Water Information System and for domain expertise that informed the anomaly taxonomy. We also thank Environment and Climate Change Canada (ECCC) for providing access to hydrometric data through the Water Survey of Canada.
The authors acknowledge the National Artificial Intelligence Research Resource (NAIRR) Pilot and the Texas Advanced Computing Center (TACC) at The University of Texas at Austin for providing computational resources under Award No. NAIRR240491.
@@
USGS streamflow observations are publicly available through the National Water Information System (NWIS) at https://waterdata.usgs.gov/nwis. Environment and Climate Change Canada hydrometric data are available through the Water Survey of Canada at https://wateroffice.ec.gc.ca/. Model weights, inference code, representative detection examples, and documentation describing the synthetic benchmark construction and Canadian station selection are available at https://huggingface.co/Ejokhan/HydroGEM.
@@
Ijaz Ul Haq: Conceptualization, Methodology, Software, Validation, Formal analysis, Investigation, Data curation, Writing - Original Draft, Visualization. Byung Suk Lee: Supervision, Resources, Writing - Review & Editing, Funding acquisition. Julia N. Perdrial: Supervision, Domain expertise, Review & Editing, Funding acquisition. David Baude: Data curation, Writing - Review & Editing.
@@ A.0.2 Tier 2: Site-Specific Standardization
$\mu_{\text{global}}=\mathbb{E}_{(s,t)\in\mathcal{D}_{\text{train}}}[\mathbf{x}_{t}^{(1)}],\quad\sigma_{\text{global}}=\text{Std}_{(s,t)\in\mathcal{D}_{\text{train}}}[\mathbf{x}_{t}^{(1)}]$ (8)
where $\mathcal{D}_{\text{train}}$ is the union of all timesteps across all training sites.
@@ A.0.5 Scale Embeddings
Design choice validation: Ablation studies (Section 5.3) confirm that removing scale embeddings substantially degrades detection performance, demonstrating their necessity for continental-scale learning.
@@ C.1 Pretraining Loss
C.2 Detection Head Training Loss
@@
where $d_{\text{RC}}$ is the absolute residual from a local power-law rating fit in log space, computed over a sliding window and averaged over timesteps.
Combined Stage 2 objective:
$\mathcal{L}_{\text{stage2}}=1.5\mathcal{L}_{\text{focal}}+\lambda_{c}(\epsilon)\mathcal{L}_{\text{corrupt}}+\lambda_{p}(\epsilon)\mathcal{L}_{\text{preserve}}+0.1\mathcal{L}_{\text{physics}}$ (38)
@@ D.0.4 Gap Statistics
  • 0% contain >>24 hours (sequences intersecting gaps >>24h are dropped)
Leakage prevention: Any sequence window intersecting a flagged timestep (outlier, physical implausibility, or gap >>24h) is excluded entirely, even if only a single hour is affected. This conservative approach eliminates potential information leakage from imputation.
@@ E.0.2 Coverage Control Algorithm
Appendix F Synthetic Test Set: Injection Formulations
This appendix provides mathematical formulations for representative anomaly types in the synthetic benchmark. Each type implements 3–4 equation variants to discourage superficial pattern matching. Additional qualitative examples, parameter tables, and reproducibility notes are available in the HydroGEM Hugging Face repository: https://huggingface.co/Ejokhan/HydroGEM.
F.1 Drift Anomaly (4 variants)
Drift represents gradual sensor calibration decay, documented by Shaughnessy et al. [67] and observed at 4.26% prevalence in labeled USGS data [61]. Four functional forms produce visually similar monotonic departures through different mechanisms.
Variant 1 – Linear:
@@ F.1 Drift Anomaly (4 variants)
Variant 4 – Polynomial:
$Q^{\prime}(t)=Q(t)+a_{Q}(t-t_{\text{start}})^{2}+b_{Q}(t-t_{\text{start}}),\quad\text{similarly for }H$ (47)
where coefficients $a,b$ are chosen to achieve target endpoint deviation $\sim U(0.1Q_{\text{mean}},0.3Q_{\text{mean}})$
F.2 Ice Backwater (3 variants)
Ice backwater occurs when ice cover elevates stage at a given discharge, a phenomenon documented in TWRI 3-A10 [31] and WSP 2175 [59]. The characteristic signature is elevated stage with suppressed or unchanged discharge.
Variant 1 – Gradual onset:
@@ F.2 Ice Backwater (3 variants)
Variant 2 – Abrupt onset with gradual recovery:
@@
Variant 3 – Periodic breakup events:
@@
where $\alpha_{\text{base}}\sim U(0.2,0.4)$, $N_{\text{events}}\sim\{2,3,4,5\}$, event amplitudes $A_{i}\sim U(-0.3,-0.1)$ (negative for breakup), widths $\sigma_{i}\sim U(1,6)$ hours
F.3 Rating Shift (3 variants)
Rating shift occurs when stream morphology changes abruptly, producing persistent bias in the stage-discharge relationship. Such shifts are documented in TWRI 3-A10 [31], and Mansanarez et al. [45] provide Bayesian methods for adjusting ratings to morphological changes at known times.
Variant 1 – Instantaneous:
$Q^{\prime}(t)=Q(t)\cdot(1+\delta),\quad t\geq t_{\text{shift}}$ (53)
where $\delta\sim U(0.15,0.55)\times\text{strength}$
Variant 2 – Transition period:
$Q^{\prime}(t)=Q(t)\cdot\left(1+\delta\cdot\frac{t-t_{\text{shift}}}{t_{\text{settle}}-t_{\text{shift}}}\right),\quad t\in[t_{\text{shift}},t_{\text{settle}}]$ (54)
where transition duration $t_{\text{settle}}-t_{\text{shift}}\sim U(6,24)$ hours
Variant 3 – Partial recovery:
$Q^{\prime}(t)=Q(t)\cdot(1+\delta\cdot r(t))$ (55)
where $r(t)$ decays from 1.0 to $r_{\text{final}}\sim U(0.3,0.7)$ over the affected window
F.4 Spike (4 variants)
Spikes are brief anomalous excursions documented at 0.13% (large) and 0.16% (small) prevalence in operational data [61]. Leigh et al. [44] include spikes in their anomaly taxonomy for water quality sensors.
Variant 1 – Electronic (instantaneous):
$Q^{\prime}(t_{\text{spike}})=Q(t_{\text{spike}})+s\cdot\sigma_{Q},\quad s\sim U(3,5)$ (56)
Variant 2 – Hydraulic (brief duration):
$Q^{\prime}(t)=Q(t)+A\cdot\exp\left[-\frac{(t-t_{\text{peak}})^{2}}{2\tau^{2}}\right],\quad\tau\sim U(1,3)\text{ hours}$ (57)
Variant 3 – Additive offset:
$Q^{\prime}(t)=Q(t)+\Delta,\quad\Delta\sim U(0.2,0.5)\times Q_{\text{segment mean}}$ (58)
Variant 4 – Bounded (capped at physical limits):
$Q^{\prime}(t)=\min(Q(t)+s\cdot\sigma_{Q},Q_{\text{max,site}})$ (59)
Appendix G Design Rationale and Parameter Selection
This appendix documents the rationale for key design decisions in SynthStream, grounded in USGS operational documentation and peer-reviewed literature.
G.1 Duration Bounds
Duration bounds for each anomaly type derive from USGS field documentation describing the temporal characteristics of sensor and hydraulic phenomena.
Table 11: Duration bounds with literature basis
Anomaly Type Duration Range Basis
Dropout 1–120 hr Telemetry outages vary from brief interruptions to multiday failures [63]
Flatline 2–144 hr Sensor freeze events persist until maintenance intervention [62]
Ice backwater 72–520 hr Ice cover persists days to weeks in northern systems [7]
Drift 96–400 hr Calibration decay develops over days to weeks [67]
Rating shift 12–288 hr Morphological changes from flood events [31, 45]
Debris effect 2–60 hr Temporary obstructions clear within hours to days [59]
G.2 Coverage Distribution
The coverage distribution (percentage of timesteps affected by anomalies) follows a trimodal design with a forbidden zone between 10% and 32%. This ensures unambiguous separation between difficulty tiers:
  • Light tier (3–9%): Sparse anomalies testing detection sensitivity
  • Moderate tier (32–44%): Substantial anomaly presence
  • Heavy tier (44–60%): Dense anomalies testing robustness
The 10–32% gap prevents ambiguous sequences that could fall into either category, enabling clean stratification of benchmark results by difficulty level.
G.3 Physical Space Injection
All anomalies are injected in physical units (ft3/s for discharge, ft for stage) rather than normalized space. This design choice ensures that injected anomalies respect the physical coupling between stage and discharge documented in rating curve theory [31]. For example, backwater anomalies elevate stage while suppressing discharge, consistent with the hydraulic relationship $Q=f(H)$ being violated by downstream control effects [59].
G.4 Equation Form Variation
Each anomaly type implements 3–4 functional forms (e.g., linear, exponential, sigmoid, polynomial for drift) that produce similar visual signatures through different mathematical mechanisms. This design prevents models from learning superficial pattern templates and requires abstraction of the underlying anomaly concept. The approach follows established practice in time series anomaly detection benchmarks [64, 56].
G.5 Single Type Sequences
Approximately 30% of sequences contain only one anomaly type, enabling unambiguous per-type evaluation. The remaining 70% contain multiple types with 40% compound overlap probability, reflecting realistic co-occurrence patterns. This split balances diagnostic clarity with operational realism.
G.6 Climate Conditional Mixing
Anomaly type probabilities are adjusted based on geographic and climatic characteristics of each station:
  • Ice backwater probability increased for high latitude stations (AK, MT, WA, MN, ND, SD, WI, MI)
  • Backwater probability increased for coastal and low elevation stations
  • Debris effect probability increased for forested watersheds
These adjustments reflect documented geographic patterns in USGS operational experience [59].